OpenStudy (anonymous):

A=(a+b)/2 solve for b please

OpenStudy (anonymous):

can you explain please

OpenStudy (anonymous):

@TAKEBACKMATH

OpenStudy (tyteen4a03):

You cannot solve for b. There are more than 1 variable. As for what @TAKEBACKMATH did, he changed the subject of the equation.

OpenStudy (anonymous):

oh okay then can you help me create my own literal equation and solve for a variable and explain how you solved for that particular variable @tyteen4a03

OpenStudy (anonymous):

no sorry, I did exactly what was asked, solved for b

OpenStudy (anonymous):

@TAKEBACKMATH is it possible to have more than 2 variables?

OpenStudy (swissgirl):

Takebackmath is correct

OpenStudy (anonymous):

thanks everyone

OpenStudy (anonymous):

@TAKEBACKMATH can you explain how you did that for me

OpenStudy (tyteen4a03):

@subbyh What is the assignment? To change subject for this equation A=(a+b)/2, here's how it's done: 2A=a+b 2A-a=b

OpenStudy (anonymous):

Hey math legend relax, I understand I need to show steps. I am a math teacher I understand

OpenStudy (anonymous):

now im confused lol the one with the explanation by @tyteen4a03 is correct??

OpenStudy (turingtest):

The goal here guys, is to guide the asker towards the answer. Please do not give the final answer unless it is unavoidable or the asker already knows it.

OpenStudy (tyteen4a03):

@subbyh What is the assignment? I think the above conversation was driving you away from the actual question.

OpenStudy (turingtest):

for instance it would be better to say: multiply by 2 subtract a then see what the asker says, no final answer needed

OpenStudy (anonymous):

it is an FLVS assignment and what was asked was to Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose. so i chose that equation i suppose i chose a rather difficult one

OpenStudy (turingtest):

not really... 1) multiply both sides by 2 2) subtract a from both sides try it, and please show your work

OpenStudy (anonymous):

i get stuck on the subtracting a part, does it matter if it is A or a

OpenStudy (tyteen4a03):

Yes. A differs from a, and capital letters are usually used to denote constants.

OpenStudy (anonymous):

2A=a+b 2A-a=b is that correct

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

thank you :)